OpenStudy (anonymous):

Multiply. (-8) x (12) A. -96 B. -4 C. 4 D. 96

OpenStudy (anonymous):

It's A :3

OpenStudy (anonymous):

I have more @DJ♬Øϟϟ¥

OpenStudy (anonymous):

positive x negative = negative positive x positive = positive negative x positive = negative negative x negative = positive

OpenStudy (anonymous):

What Is the sign of the product? (0.5)(-0.5)(3.2)(-7.9)(-2.4) A.Positive B.Negative

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

OpenStudy (anonymous):

@hippo4

OpenStudy (anonymous):

@alexwright @Ocabre

OpenStudy (anonymous):

-15.17 So, it's negative

OpenStudy (anonymous):

Yes I agree ^^

OpenStudy (anonymous):

FInd The mean of this set of numbers 5, -11, 7, 12, -1, 0 a. 2 b. -6 c. 6 D. -2

OpenStudy (anonymous):

@DJ♬Øϟϟ¥ @alexwright

OpenStudy (anonymous):

The answer is 2

OpenStudy (anonymous):

yep ^^ @DJ♬Øϟϟ¥

OpenStudy (anonymous):

Simplify 12-16/ (-4) a. 16 b.8 c. 1 d. -1

OpenStudy (anonymous):

@alexwright @DJ♬Øϟϟ¥

OpenStudy (anonymous):

It's C :)

OpenStudy (anonymous):

wait how did you get C when I put it in my calculator I got 16? @DJ♬Øϟϟ¥

OpenStudy (anonymous):

Evalulate the expression when p+ -16 and q+ -2 a.2 b. -8 c. -2 d. 8

OpenStudy (anonymous):

12-16= -4 -4/-4= 1 (negative divided by a negative is a positive) :)

OpenStudy (anonymous):

im not saying you got it wrong I wanna know where I messed up for the future?

OpenStudy (anonymous):

oh okay stupid calculator thanks :)

OpenStudy (anonymous):

Wait i typed that wrong p= -16 and q=-2 @alexwright @DJ♬Øϟϟ¥

OpenStudy (anonymous):

XD You're welcome. And if you put -4/-4, it will change the division sign into a minus sign. :)

OpenStudy (anonymous):

Wait, you didn't give me the expression :P

OpenStudy (anonymous):

p over 4q

OpenStudy (anonymous):

-16/4*-2 The answer is 2 :)

OpenStudy (anonymous):

Which value of x makes this equation true? -8 + x=0 a. -8 b. 8 c. no value of x will make it true d. 0

OpenStudy (anonymous):

@dj

OpenStudy (anonymous):

It's B :P

OpenStudy (anonymous):

Round the number to the hundreaths place 12,863. 14846 a. 12,900 b. 12,800 c. 12,863.15 d. 12,863.14

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

OpenStudy (anonymous):

It's C

OpenStudy (anonymous):

Estimate the value of this expression by rounding each number to the nearest whole unit (1.74)(9.95)(3.46) a. 60 b. 80 c. 54 d. 72

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

OpenStudy (anonymous):

2*10*3=60 So, A

OpenStudy (anonymous):

Find the value of x that makes the equation true 7x=-56 a.-392 b. -8 c. -63 d.-49

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

OpenStudy (anonymous):

It's B

OpenStudy (anonymous):

identify which property of equality is used to transform 1 to equation 2. equation 1: 3x=12 equation 2: x=4 a.addition property to equality b/division property to equality c. multiplication property to equality d. subtraction property to equality

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

OpenStudy (anonymous):

I'm going with C.

OpenStudy (anonymous):

solve the equation for x 9x= -5y a. x= -45y b. x=-5y over 9 c. x= -14y d. x= - 9 over 5y

OpenStudy (anonymous):

@DJ♬Øϟϟ¥

OpenStudy (anonymous):

It's B